UT 20150507 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson
Observer AO: DMiller (LBTO)
Telescope Operator: SAllenson
LBTO Support: TSargent

Plan for the night

- Start with AO-Com#3, setup the system within its limitations

- Do tasks from ComPlan depending on how it goes

Summary

  • Startup AO successfull

  • Checked and corrected AO stage transform

  • Photometric ZPs N30-camera

Details

All times are UT
Files are luci2.20150507.NNNN.fits

Had an MOS-error in the afternoon (see attached log), DThompson was able to recover!

02:10 Sunset, mostly clear!!!, T = 3deg. Hum = 40%, Wind ~ 10m/s
For the moment being, we should avoid observing towards SW

02:34 Steve sends preset, checks pointing and collimates on A45_R8_0031
collimation works moderately well, seeing is about 1.2"

Super, star lands already on the Pyramid without a need to search for it

02:50 DMiller is taking over, he tries to close the loop, unfortunately his GUIs are frozen, reboot OBS3

03:12 Steve restarts AOS, Doug has problems with the arbitrators, which crash frequently, contacts Alfio

03:20 Wind is increasing, about 15m/s now, seeing improves 0.7"

0065 star @ 1017 1029 on N3.75
0066 star @ 1021 1026 on N3.75 w/loops closed
0067 star @ 882 792 on N30 w/loops closed FWHM=7.2px, offset 2.13 3.48 rel DETXY
0068 star @ 1030 1022 on N30
Update /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf (saved copy first):
RIGHTPOINTINGORIGIN 2.13 3.48
Stop and restart PCS, re-send same preset
0069 star @ 1080 1084
Looks like the values should be in mm, not arcsec
RIGHTPOINTINGORIGIN 1.28 2.09
Stop and restart PCS, re-send the same preset
0070 star @ 995 1022 AO loops still open

0072 star @ 980 1007 AO loops closed
Tried PA+180 but no guide star. AOref is visible on the pyramid though

Moving to another star. guide stars available at -45 and +135 PAs

Note on RTD the ends of the filenames are cut off at right. Can this be corrected? Dragging the window wider does not help. E-mail sent to Ingo and Jochen.

Apparently we still have a restriction on available rotator angles that we can use and get the AO loops closed reasonably stably (between 130 and 240 degrees).

###Check AO stage transform currently in place

0073 00xx 3 x 3.0s BrG N30 BS9158_AOtrans.xml
Aborted...script requests N30 camera but UI thinks it is in N3.75,changed script

0074 0098 BS9166_AOtrans.xml @PA=0, rotation is 1.68 degrees clockwise on detector
0099 image of the N30 LS neu 1 longslit (rotated 0.119 degrees clockwise on detector)

pattern must be offset 1.56 degrees from rotation matrix already in place (-0.3 degrees)

0100 01xx BS9166_AOtrans.xml @PA=0
rotation offset of 1.56 degrees from existing transform is applied

0100-0103 reposition field stop on detector

0104 BS9166_AOtrans.xml @ PA=0
new transform rotated by 1.56 degrees in place.
I moved the FM4 too far (used the wrong scale for N30)...corrected

0123 0124 BS9166_AOtrans.xml
AO resume failed before third position

0125 0127 BS9166_AOtrans.xml
AO resume failed again

Made a new script that only does 2 positions +/-5" in Y
0128 01xx BS9166_AOtrans2.xml
Rotation was 3.67 degrees...went the wrong way

0139 0146 BS9177_AOtrans1.xml @PA=70
two positions off the top of the detector
modified positions, reduced offset gtom 9 to 6 arcsec

0147 0157 BS9177_AOtrans2.xml @PA=70
pause after first exposure to check...source is still ~3" above center
rotation seems to be good now. Difference between offsets and slit is about 0.07 deg.

Its photometric outside, doing photo ZPs

0158 0162 ZP_BS9177_N30_long.xml (modified to request N3.75(N30))
UI appears to be frozen, multiple calls to camera needed to be fixed in the script

0163 test image to make sure we were still in the N30 camera

0164 0223 ZP_BS9177_N30_long.xml (modified to request N3.75(N30) always)
Stopped at the end of the Jhigh dataset

0224 0225 Reconfigure to N3.75

0226 0250 ZP_BS9177_N3.75.xml

Guiding and wavefront sensing lost...done for the night.

12:25 Sunrise, Steve is closing up

Things that need doing:
- newpoint.cl under /lbt/lbttools/Observe was changed to allow ACE_ACTIVE presets to be sent (this is hard-coded).
It should be checked into SVN.

- RTD filename list at right is truncated. Ingo emailed to see if this can be fixed.

- M3 position or whatever is causing the AO pointing to be a function of the rotator angle needs to be corrected.

-- %USERSIG{JochenHeidt - 2015-05-07}%
Topic revision: r4 - 19 May 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback